OpenStudy (anonymous):

What is the equivalent to (mn)3?

OpenStudy (bibby):

\((ab)^m=a^mb^m\)

OpenStudy (anonymous):

I was just asked this, but I have never been taught it before. This is completely new to me. This is not the only one though.

OpenStudy (anonymous):

@bibby thank you, but that isn't one of the answers.... :/

OpenStudy (bibby):

that isn't an answer. it's a generalized approach to problems like these a, b, and m are numbers for example: \( (ab)^{-2}=a^{-2}*b^{-2}\)

OpenStudy (anonymous):

Ahhhhh.... OK then. I am in 8th grade and my teacher has never actually taught anything like this and now it is on the Segment 1 exam.... If you want to see the answers they asked, here they are: 3mn; m3n; mn3; m3n3. And, if you don't mind, please explain to me how you got the answer.

OpenStudy (bibby):

my point is the exponent carries on to the individual terms it's a property of exponents if \((ab)^2=a^2b^2\) then apply that same logic to your problem

OpenStudy (anonymous):

Thank you!